# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ B.3.2.  Regret Upper Bound
$\lesssim\frac{m\log M}{\eta}+\eta Tm^{2}M^{3}$
Setting $\eta=\frac{1}{M}\sqrt{\frac{\log M}{mMT}}$, we get $\mathsf{REG}\leq O(M\sqrt{m^{3}MT\log M})\,.$  [^10] Recall that the analysis assumes that $\eta\leq\nicefrac{{1}}{{M}}$. As $m\geq 1,T\geq 1$ and $\log M\leq M$, the chosen $\eta=\frac{1}{M}\sqrt{\frac{\log M}{mMT}}\leq\frac{1}{M}.$, [^11] The upper bound on $\mathsf{REG}$ has a linear dependence on $M$ which is unavoidable as the reward in each round is in $[0,M]$. If normalized regret (Zhang et al., 2024) is considered instead, i.e., the rewards are scaled such that in each round it lies in $[0,1]$, the dependence on $M$ is logarithmic (or sublinear in the bandit setting).
B.3.3.  Time Complexity
In the full information setting, the running time bottleneck is Eq. 12. Thus, Algorithm 1 runs in $|E|=O(mM^{2})$ time per round. The bottleneck of running Algorithm 1 under the bandit feedback is the efficient computation of the marginal distribution, $p^{t}(e)$. Note that $p^{t}(e)$ has a sum-product structure as follows:
$p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\prod_{e^{\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime})\,.$
Hence, by the message passing algorithm from Wainwright and Jordan (2008, Section 2.5.1), we can compute the marginal probabilities (and hence run Algorithm 1) in $O(M|E|)=O(mM^{3})$ time per round.
@@ B.3.3.  Time Complexity
For $m=1$, let $K=M$ be an even integer. Let $\mathbf{v}=[1,\dots,1,v,\dots,v]$. Here, the first $\frac{M}{2}$ entries are 1 and the remaining $\frac{M}{2}$ entries are $v$. Let $v=1-\delta$. Define $\delta^{\prime}=\frac{\delta}{2M}$. Consider two scenarios:
@@
$\bm{\beta}_{-}^{t}=\begin{cases}[1-\delta^{\prime},\dots,1-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}+\delta,\\ [\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}-\delta,\\ \end{cases}$
@@
$\bm{\beta}_{-}^{t}=\begin{cases}[1-\delta^{\prime},\dots,1-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}-\delta,\\ [\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}+\delta,\\ \end{cases}$
for some $\delta\in(0,1/4)$ that would be determined shortly. Assume the randomness used in different rounds are independent.
Let $P$ and $Q$ be the distribution of $[\bm{\beta}_{-}^{t}]_{t\in[T]}$ for scenario 1 and 2 respectively. Then, for $\delta\in(0,\frac{1}{4})$,
@@
Hence, no strategy of form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$ is allocated any unit in when the competing bids are $[1-\delta^{\prime},\dots,1-\delta^{\prime}]$. The smallest bid value for $\frac{M}{2}+1\leq q\leq M$ is for $q=M$. Note that, $w_{M}=\frac{1+v}{2}>\frac{1+v}{2}-\delta^{\prime}$. Hence, $(w_{M},M)=(\frac{1+v}{2},M)=(1-\frac{\delta}{2},M)$ is allocated all the units when the competing bids are $[\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}]$ and, by definition, weakly dominates all the strategies of the form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$. Hence, we have two undominated (for the constructed competing bids) strategies: $\mathbf{b}_{1}=(1,\frac{M}{2})$ and $\mathbf{b}_{2}=(1-\frac{\delta}{2},M)$.
Now, we compute the expected value obtained by $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ for scenarios $P$ and $Q$.
@@
So, $\mathbf{b}_{1}$ (resp. $\mathbf{b}_{2}$) is optimal for scenario $P$ (resp. scenario $Q$). Now, consider the distribution $\frac{P+Q}{2}$, i.e., $\textsc{Bern}(0.5)$.
$\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]=\frac{M}{2}~{}~{}\text{ and }~{}~{}$ $\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]=\frac{M}{2}\left(\frac{1+v}{2}\right)\leq\frac{M}{2}$
@@
$\geq\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]-2\max_{b^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
@@
(13)${{\geq}}\frac{M}{2}+\frac{M}{2}\left(\frac{1}{2}+\delta\right)(2-\delta)-2\cdot\frac{M}{2}=\frac{M}{2}\left(\frac{3\delta}{2}-\delta^{2}\right)\geq\frac{5M\delta}{8},$
@@
$\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]\geq\frac{5MT\delta}{16}\cdot(1-\text{TV}(P,Q))\geq\frac{5MT\delta}{32}\exp{\left(-16T\delta^{2}\right)}$
Setting $\delta=\frac{1}{4\sqrt{2T}}$, we get $\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]\geq\frac{M\sqrt{T}}{36\sqrt{e}}$. Hence, $\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]=\Omega(M\sqrt{T})$.
@@ Appendix C Omitted Proofs from Section 5
C.1.  Proof of Theorem 5.3
To prove that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}\leq 2$ for any $m\in\mathbb{N}$, recall from the proof sketch that we first defined the following metric for any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (14)
Let $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ where $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ [^12] We assumed $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ without loss of generality. The proof also follows if we considered $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{k}^{*},q_{k}^{*})\rangle$ for some $k\in[m]$ instead.Define $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$. Suppose $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}^{*}$ units in round $t$. For any $j\in[m]$, let $T_{j}$ be the set of rounds in which the least winning bid is $b_{j}^{*}$, i.e.,
$T_{j}=\big{\{}t\in[T]:Q_{j-1}^{*}<r_{t}^{*}\leq Q_{j}^{*}\big{\}}\,.$ (15)
$T_{j,0}=\Big{\{}t\in[T_{j}]:r_{t}^{*}<Q_{j}^{*}\Big{\}},\qquad T_{j,1}=\Big{\{}t\in[T_{j}]:r_{t}^{*}=Q_{j}^{*}\Big{\}}\,.$ (16)
For any $j\in[m]$, let
$\widehat{Q}_{j}=\begin{cases}\max\{r_{t}^{*}:t\in T_{j,0}\},~{}&\text{ if }T_{j,0}\neq\emptyset\\ Q_{j}^{*}~{}&\text{ if }T_{j,0}=\emptyset\,.\end{cases}$
Recall that any strategy in $\mathscr{U}_{m}^{\star}$ with $m$ bid-quantity pairs takes the form of $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},\forall\ell\in[m]$. The strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ also have the same structure but as an important difference, for any $\ell\in[m]$, we enforce $Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\}$. Observe that for any $\ell\in[m-1]$,
$Q_{\ell}\leq Q_{\ell}^{*}<\widehat{Q}_{\ell+1}\leq Q_{\ell+1},$
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (18)
Step 2. Value Decomposition for $m$-uniform Strategies. Let $V_{j,k}$ be the time-average value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in the set of rounds in $T_{j,k}$ (as defined in Eq. 16). Formally, $\forall j\in[m],k\in\{0,1\}$,
Define ${W}_{k}=\sum_{j=1}^{k}v_{j}$. Note that $V_{j,1}={W}_{Q_{j}^{*}}$ because for any $t\in T_{j,1}$, we have $r_{t}^{*}=Q_{j}^{*}$. For any $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}}$, and $Q_{\ell}\in\{Q_{\ell}^{*},\widehat{Q}_{\ell}\}$ for any $\ell\in[m]$, define $N^{*},\widehat{N}\subseteq[m]$ as follows:
$N^{*}=\{j:Q_{j}=Q_{j}^{*},j\in[m]\}\quad\text{ and }\quad\widehat{N}=\{j:Q_{j}=\widehat{Q}_{j},j\in[m]\}\,.$
We now present a crucial result that expresses the value obtained by a $m$-uniform bidding strategy as a function of the value obtained by $m$ 1-uniform strategies.
Lemma C.1.
where we recall that $Q_{\ell}=\sum_{j=1}^{\ell}q_{j},\forall\ell\in[m]$.
A key consequence of Lemma C.1 is that knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$ is sufficient to compute the value obtained any $m$-uniform safe bidding strategy. By Lemma C.1, for any round $t\in T_{j}$, and $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, we get
$\implies\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $\geq\sum_{j\in N^{*}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})+\sum_{j\in\widehat{N}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,.$ (20)
Step 3. Allocation Lower Bounds for 1-uniform Strategies. For any $j\in[m]$, we now invoke the following lemma to establish lower bound on $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$
Lemma C.2.
  • (a)
    if $Q_{j}=\widehat{Q}_{j}$ (i.e., $j\in\widehat{N}$), we have
  • (b)
    If $Q_{j}=Q_{j}^{*}$ (i.e., $j\in N^{*}$), we have
Note that the right hand side of Eq. 20 depends only on the choice of the partitions $N^{*}$ and $\widehat{N}$. Substituting the lower bound from Lemma C.2 in Eq. 20, we establish that,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (18)${{\leq}}\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}$
$\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{(N^{*},\widehat{N})}\Big{\{}\sum_{j\in N^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}}\,.$ (23)
Let $(N_{0}^{*},\widehat{N}_{0})=\mathop{\mathrm{argmax}}_{(N^{*},\widehat{N})}\Big{\{}\sum_{j\in N^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}$. Then,
$=\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}\,.$ (24)
Consider a partition of $[m]$ that differs from the maximizing partition $(N_{0}^{*},\widehat{N}_{0})$ by exactly one element, i.e., for any $a\in\widehat{N}_{0}$, consider the following partition: $(N_{0}^{*}\cup\{a\},\widehat{N}_{0}\setminus\{a\})$. By definition,
$\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$ $\geq\sum_{j\in N_{0}^{*}\cup\{a\}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}\setminus\{a\}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$
Now, for any $b\in N^{*}_{0}$, consider the following partition: $(N_{0}^{*}\setminus\{b\},\widehat{N}_{0}\cup\{b\})$. By definition,
$\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$ $\geq\sum_{j\in N_{0}^{*}\setminus\{b\}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}\cup\{b\}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$
Plugging in the values,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (C.1)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
(25)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}2V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
(26)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}2{W}_{Q_{j}^{*}}|T_{j,1}|-{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}2V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
$=\frac{2\Big{\{}\sum_{j\in N_{0}^{*}}{W}_{Q_{j}^{*}}|T_{j,1}|+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}-\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
$=2-\theta_{\mathcal{H}_{-}}\,,$
where
$0<\theta_{\mathcal{H}_{-}}$ $=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
(25)${{\leq}}\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}}$
$=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j=1}^{m}{W}_{Q_{j}^{*}}|T_{j,1}|}\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{j\in[m]}\frac{{W}_{\widehat{Q}_{j}}}{{W}_{Q_{j}^{*}}}\leq 1\,,$
and $(a)$ follows from 1. Finally,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}=\max_{\mathcal{H}_{-}}\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq 2-\min_{\mathcal{H}_{-}}\theta_{\mathcal{H}_{-}}=:2-\theta,\quad\text{where}\quad\theta=\min_{\mathcal{H}_{-}}\theta_{\mathcal{H}_{-}}\in(0,1]\,.$
Fact 1 (Williamson and Shmoys (2011, pp. 25)).
For positive numbers $a_{1},\dots,a_{k}$ and $b_{1},\dots,b_{k}$,
$\frac{\sum_{j=1}^{k}a_{j}}{\sum_{j=1}^{k}b_{j}}\leq\max_{j\in[m]}\frac{a_{j}}{b_{j}}\,.$
C.2.  Proof of Lemma C.1
We state and prove a stronger version of the result in Lemma C.1. Formally,
@@
Lemma C.3 states a similar result as Lemma C.1, except a key difference that the bidding strategies are not necessarily safe.
C.2.1.  Proof of Lemma C.3
@@
Case II. $r>Q_{m}$. As $r>Q_{m}$, $b_{m+1}$ is the least winning bid which implies $b_{m+1}\geq\bm{\beta}_{-}^{-(r)}$, where we recall that $\bm{\beta}_{-}^{-(r)}$ is the $r^{th}$ smallest competing bid in $\bm{\beta}_{-}$. So, $(b_{m+1},Q_{m+1})$ is allocated at least $r$ units which implies $V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})\geq V(\mathbf{b};\bm{\beta}_{-})$. By C.4, we also have $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})$. Hence,
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\max\Big{\{}\max_{\ell\in[m]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big{\}}$

### v4

@@ Appendix B Omitted Proofs From Section 4
$\lesssim\frac{m\log M}{\eta}+\eta Tm^{2}M^{4}\,.$
Setting $\eta=\frac{1}{M^{2}}\sqrt{\frac{\log M}{mT}}$, we get $\mathsf{REG}\leq O(M^{2}\sqrt{m^{3}T\log M})$.
Time Complexity. The bottleneck of running Algorithm 1 under the bandit feedback is the efficient computation of the marginal distribution, $p^{t}(e)$. We claim that this can be done in $O(|E|)=O(mM^{2})$ time per-round.
We begin by defining a quantity analogous to $\Gamma^{t-1}(\cdot)$. Define $\widetilde{\Gamma}^{t-1}(s)=1$ and $\widetilde{\Gamma}^{t-1}(\cdot)$ is computed recursively for every node $u\in V$ as follows:
$\widetilde{\Gamma}^{t-1}(u)=\sum_{v:v\to u\in E}\widetilde{\Gamma}^{t-1}(v)\cdot\varphi^{t-1}(v\to u)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(v\to u))\,.$ (32)
Claim B.3.
For any node $u$ in the graph, let $\widetilde{\mathscr{P}}(u)$ be the set of paths starting at $s$ and terminating at $u$. Then,
$\widetilde{\Gamma}^{t-1}(u)=\sum_{\mathfrak{p}\in\widetilde{\mathscr{P}}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))$
Proof.
The proof follows the same structure as B.1, but instead of proceeding bottom-up through the nodes, we use top-down induction. ∎
Let $e=u\to v$. Recall that
[…]
$=\varphi^{t}(e)\left(\sum_{\mathfrak{p}\in\widetilde{\mathscr{P}}(u)}\prod_{e^{\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime})\right)\left(\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e^{\prime\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime\prime})\right)$
[…]
$\quad\times\left(\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e^{\prime\prime}\in\mathfrak{p}}\varphi^{t-1}(e^{\prime\prime})\exp(\eta\widehat{\textsf{w}}^{t-1}(e^{\prime\prime}))\right)\,.$
By B.1, B.3, Eq. 30 and the fact that $\Gamma^{t-1}(d)=1$, we get
$p^{t}(e)$ $=\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}\cdot\frac{\Gamma^{t-1}(u)}{\Gamma^{t-1}(s)}\cdot\frac{1}{\Gamma^{t-1}(v)}\cdot\widetilde{\Gamma}^{t-1}(u)\cdot\Gamma^{t-1}(v)$
$=\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))\cdot\frac{\widetilde{\Gamma}^{t-1}(u)\Gamma^{t-1}(v)}{\Gamma^{t-1}(s)}$
Observe that $\Gamma^{t-1}(\cdot)$ (resp. $\widetilde{\Gamma}^{t-1}(\cdot)$) can be computed for all nodes recursively using Eq. 31 (resp. Eq. 32) in $O(|E|)$ time. Hence, $p^{t}(e)$ can be computed for all edges $e\in E$ in $O(|E|)=O(mM^{2})$ time in any round.
@@
For $m=1$, let $K=M$ be an even integer. Let
$\mathbf{v}=[1,\dots,1,v,\dots,v]\,.$ (33)
Here, the first $\frac{M}{2}$ entries are 1 and the remaining $\frac{M}{2}$ entries are $v$. Let $v=1-\delta$. Define $\delta^{\prime}=\frac{\delta}{2M}$,
$\bm{\beta}_{-}^{\clubsuit}=[1-\delta^{\prime},\dots,1-\delta^{\prime}]\quad\text{and}\quad\bm{\beta}_{-}^{\diamondsuit}=\Big[\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}\Big]$ (34)
Consider two scenarios:
@@
[…]
@@
[…]
for some $\delta\in(0,1/4)$ to be determined shortly. Assume the randomness used in different rounds are independent.
Let $P$ and $Q$ be the distribution of $\bm{\beta}_{-}^{t}]_{t\in[T]}$ for scenario 1 and 2 respectively. Then, for $\delta\in(0,\frac{1}{4})$,
@@
Hence, no strategy of form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$ is allocated any unit in when the competing bid profile is $\bm{\beta}_{-}^{\clubsuit}$. The smallest bid value for $\frac{M}{2}+1\leq q\leq M$ is for $q=M$. Note that, $w_{M}=\frac{1+v}{2}>\frac{1+v}{2}-\delta^{\prime}$. Hence, $(w_{M},M)=(\frac{1+v}{2},M)=(1-\frac{\delta}{2},M)$ is allocated all the units when the competing bid profile is $\bm{\beta}_{-}^{\diamondsuit}$ and, by definition, weakly dominates all the strategies of the form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$. Hence, we have two undominated (for the constructed competing bids) strategies: $\mathbf{b}_{1}=(1,\frac{M}{2})$ and $\mathbf{b}_{2}=(1-\frac{\delta}{2},M)$.
Now, we compute the expected value obtained by $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ for scenarios 1 and 2.
@@
So, $\mathbf{b}_{1}$ (resp. $\mathbf{b}_{2}$) is optimal for scenario 1 (resp. scenario 2). Now, consider the distribution $\frac{P+Q}{2}$, i.e., $\textsc{Bern}(0.5)$.
[…]
@@
$\geq\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]-2\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
@@
[…]
@@
$\mathbb{E}_{\frac{P+Q}{2}}[\textsf{REG}]\geq\frac{5MT\delta}{16}\cdot(1-\text{TV}(P,Q))\geq\frac{5MT\delta}{32}\exp{\left(-16T\delta^{2}\right)}$
Setting $\delta=\frac{1}{4\sqrt{2T}}$, we get $\mathbb{E}_{\frac{P+Q}{2}}[\textsf{REG}]\geq\frac{M\sqrt{T}}{36\sqrt{e}}$. Hence, $\mathbb{E}_{\frac{P+Q}{2}}[\textsf{REG}]=\Omega(M\sqrt{T})$.
@@ Appendix C Omitted Proofs from Section 5
C.1. Proof of Lemma 5.6
We state and prove a stronger version of the result in Lemma 5.6. Formally,
@@
Lemma C.1 states a similar result as Lemma 5.6, except a key difference that the bidding strategies are not necessarily safe.
C.1.1. Proof of Lemma C.1
@@
Case II. $r>Q_{m}$. As $r>Q_{m}$, $b_{m+1}$ is the least winning bid which implies $b_{m+1}\geq\bm{\beta}_{-}^{-(r)}$, where we recall that $\bm{\beta}_{-}^{-(r)}$ is the $r^{th}$ smallest competing bid in $\bm{\beta}_{-}$. So, $(b_{m+1},Q_{m+1})$ is allocated at least $r$ units which implies $V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})\geq V(\mathbf{b};\bm{\beta}_{-})$. By C.2, we also have $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})$. Hence,
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\max\Big\{\max_{\ell\in[m]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big\}$